# A calibrated model for N-body dynamical friction acting on supermassive black holes

### v1

@@ Abstract
pubyear: 2023pagerange: A calibrated model for N-body dynamical friction acting on supermassive black holesA
@@ 1.2 Compensating for unresolved dynamical friction
Finally, the method of Pfister et al. (2019) is similar to those described above, but takes into account the contribution of fast-moving particles to dynamical friction, which may be important in enhancing dynamical friction in cases where the density distribution of the background particles becomes shallow (Antonini & Merritt, 2012; Dosopoulou & Antonini, 2017), such as in central density cores (de Blok, 2010). However, we also note that the “standard” Chandrasekhar treatment is known to fail in reproducing the effect of stalling in cored density distributions (Read et al., 2006; Banik & van den Bosch, 2021).
@@ 1.3 Dynamical friction from $N$-body particle distributions
Using their method, the dynamical friction force can be computed straightforwardly with direct summation, or with tree methods by treating a distant node as a single particle. Manifest momentum conservation can also be introduced by applying an opposite force to each particle that contributes dynamical friction to the black hole. Nevertheless, for relatively well-resolved cases, such an approach, in effect, double-counts the dynamical friction because a fraction of the force is already resolved in the simulation and taken into account in the force computation. Ma et al. (2023) acknowledge this shortcoming and point out that this can be corrected in the future by some mass-resolution-dependent weight function.
@@ 2 Methods
Table 1: The black hole to dark matter mass ratio, the number of particles used, the number of particles within the original orbital radius of the black hole and the employed gravitational softening for our primary set of simulations carried out with GADGET-4.
@@
5000 $5\times 10^{7}$ $2\times 10^{6}$ 0.20
@@ 2.3 Predictions from simulations with KETJU
In order to make sure that the Chandrasekhar description works for our set-up, we also run a high-resolution $N$-body simulation using the KETJU code (Rantala et al., 2017; Mannerkoski et al., 2023). The essence of the GADGET-4-based KETJU code is to add a regularized region around each massive black hole in the simulations in which the gravitational dynamics is accurately integrated without gravitational softening. Throughout this work we use the regularized region size of $2.8\,\epsilon_{g}$ in radius, resulting in region sizes up to a few hundreds of pc, ensuring that all interactions with the black hole in the simulations are always non-softened. The interactions between dark matter particles in the region remain softened. This allows us to capture the dynamical friction below the softening length, while also accounting for deviations from the Chandrasekhar approximation (e.g. like the assumption of hyperbolic orbits). Moreover, since a number of previous methods used to include unresolved dynamical friction in simulations rely on the Chandrasekhar description, we aim to learn about any potential shortcomings of these methods through this test.
In KETJU, the integration around black holes (the MSTAR integrator, Rantala et al. 2020; Mannerkoski et al. 2023) relies on algorithmic regularization techniques (Mikkola & Tanikawa, 1999; Mikkola & Merritt, 2006), a minimum spanning tree coordinate system and the Gragg-Bulirsch-Stoer (GBS) extrapolation method (Gragg, 1965; Bulirsch & Stoer, 1966) to achieve an extremely high integration accuracy. The efficient two-fold parallelization of MSTAR allows for integrating up to $\sim 5000$ particles in a single regularized region. In order to properly resolve dynamical friction, KETJU requires a mass ratio of $M_{\rm BH}/M_{\rm DM}\gtrsim 100$ between the black hole and the dark matter particles (Rantala et al., 2017). KETJU has been successfully applied for simulating massive black hole and black hole binary dynamics in both isolated galaxy mergers (e.g. Rantala et al. 2018) and cosmological zoom-in simulations (Mannerkoski et al., 2022). Note that in this work we employ only Newtonian gravity and do not include post-Newtonian terms in the equations of motion close to the black hole, something that is in principle available in KETJU, as the dark matter mass elements do not represent physical particles. For the user-given accuracy parameters of KETJU, we adopt a GBS tolerance of $\epsilon_{\mathrm{GBS}}=10^{-7}$ and an integration end-time tolerance of $\epsilon_{\mathrm{t}}=10^{-2}$.
@@ 3 Performance of existing methods
The result of applying the Chen et al. (2022) formula can be seen in the middle panel of Fig. 2, where the black holes now seem to sink, on average, slower than the prediction of KETJU by up to 1 Gyr, and there appears to be a slight resolution-dependent gradient in the sinking time. This may suggest that $b_{\rm max}=6\,\epsilon_{g}$ in the Coulomb logarithm is too large of a dynamical friction correction, particularly in cases where dynamical friction is to a large part resolved.
@@ 4 A calibrated dynamical friction model
We note that the order of magnitude of the second-order coefficients is determined primarily by the mass of the background particles. The coefficient in Eq. (11) acts opposite to the direction of the dynamical friction force and accounts for the velocity dispersion of the slow-moving background particles and the typical velocity of the fast-moving particles.
@@
Through explicit evaluation of Eqns. (10), (11) and (12), we have determined that for an isotropic Hernquist halo the second-order diffusion term acting parallel to the direction of motion of the black hole, ${\bf a}_{\rm noise,\parallel}$, approximately cancels out the additional cooling term, ${\bf a}_{\rm df,m_{*}}$, and only dominates near the centre of the halo. The second-order perpendicular term, ${\bf a}_{\rm noise,\perp}$, is approximately equal in magnitude to the parallel term, suggesting that Eq. (14) is a good approximation to the magnitude of the noise terms. However, we also found that the Maxwellian approximation typically overestimates the stochastic term in a Hernquist halo by a factor of $\sim 3$, while the numerical evaluation of local $\sigma_{\rm 3D}/v$ is extremely noisy and can lead to order-of-magnitude changes in the black hole acceleration. Moreover, it is unclear how one would account for the second-order terms perpendicular to the black hole motion, as we cannot at any given point determine the net direction of the stochastic force.
@@ 4.3 Probing different values of the softening
$\eta_{\rm optimal}=\rm max[0.46~{}({\epsilon_{g}}/{\rm kpc})^{0.17}~{}10^{~{}0.11\ln^{2}(\epsilon_{g}/{\rm kpc})}~{},~{}1.0],$ (16)
where we fixed the maximum value of $\eta$ to 1. The relation $b_{\rm max}=\epsilon_{g}$ is shown with the grey dashed line. We note that the definition of $b_{\rm min}$ of Tremmel et al. (2015) and Chen et al. (2022) is different to ours, so our relation would not hold in that case; however, with $b_{\rm min}=GM/v^{2}_{\rm BH}$, and through an approximation $v^{2}_{\rm BH}\approx\frac{1}{3}\sigma_{\rm 3D}^{2}$, we can deduce that an optimal value is $\eta\approx 2$ for Tremmel et al. (2015) and Chen et al. (2022). We show the optimal value of $\eta$ with this definition of $b_{\rm min}$ in Appendix A.
@@
$\eta_{\rm optimal}=\rm max[103.7~{}({\epsilon_{g}/a_{\rm HQ}})^{1.3}~{}10^{~{}0.11\ln^{2}(\epsilon_{g}/a_{\rm HQ})}~{},~{}1.0].$ (17)
@@ 4.6 A treatment for low-mass black holes
Figure 9: Left: the effect of increasing the dynamical mass of black holes on their orbital trajectories. In black we show the analytical prediction for the orbit of a $10^{8}\,{\rm M}_{\odot}$ black hole in a $10^{13}\,{\rm M}_{\odot}$ halo, starting on a circular orbit at 10 kpc. The initial orbit with $M_{\rm BH}/M_{\rm DM}=1$ is shown with a thin dark blue line. The dynamical mass of the black hole is then scaled by a factor $k$ (see legend), resulting in faster black hole sinking. No dynamical friction correction is applied. The thick blue line shows the result of applying our model coupled with the stochasticity correction via a rescaling of the dynamical mass. Right: same as the left panel, but for a $10^{7}\,{\rm M}_{\odot}$ black hole starting at 5 kpc. The black hole trajectories have been smoothed over 0.1 Gyr for clarity.
@@
where $M_{\rm DM}$ is the background particle mass, which brings the noise correction to the right order of magnitude (see Eqns. 11 and 12). Since particle scattering below $\eta\epsilon_{g}$ is unresolved, only impact parameters above $\eta\epsilon_{g}$ contribute, giving the Coulomb logarithm $\ln\Lambda=\frac{R}{\eta\epsilon_{g}}$. We note that this is different to applying the term ${\bf a_{\rm df,m_{*}}}$ as in the work of Damiano et al. (2024) – their additional “cooling” force accounts for unresolved dynamical friction from the mass of background particles (assuming their masses are the true masses of background stars and dark matter), while Eq. (20) attempts to negate the resolved heating of the black hole by massive background particles.
@@
The result of the application of this correction through a rescaling of the dynamical mass is shown with a thick blue line in Fig 9. In neither case does the correction fully reproduce the true trajectory of the black hole, however, in both cases the black holes sink to the centre of the galaxy without the need for a substantial increase of their dynamical mass.
@@ 4.7 Black holes in the regime $1\leq M_{\rm BH}/M_{\rm DM}\leq 5$
Figure 10: The results of applying our dynamical friction correction, coupled with the correction for stochasticity, to a $10^{7}\,{\rm M}_{\odot}$ black hole in a $10^{13}\,{\rm M}_{\odot}$ halo starting at 4 kpc. Darker lines show median black hole trajectories for different $M_{\rm BH}/M_{\rm DM}$ ratios as the halo mass resolution changes. The arrows indicate the earliest and the fastest sinking black holes from the sample of 6 initial conditions used to compute the median trajectories. The lighter lines show examples of trajectories where only a dynamical friction correction is applied, i.e. no stochasticity correction. The black hole trajectories have been smoothed over 0.15 Gyr for clarity.
We now explore how our stochasticity correction behaves in the regime $1\leq M_{\rm BH}/M_{\rm DM}\leq 5$. This is shown in Fig. 10, where for a fixed halo and black hole mass, we gradually increase the mass resolution such that $M_{\rm BH}/M_{\rm DM}$ varies from 1 to 5. The dark lines show the median trajectory of 6 black hole seeds for each mass ratio. We apply the stochasticity correction via Eq. (20). The correction clearly tends, on average, to sink the black holes too fast, with the effect being strongest for lower $M_{\rm BH}/M_{\rm DM}$. The arrows show the earliest and latest sinking times of the black holes, where the mass ratio appears to set the earliest black hole sinking time, while the maximum sinking time does not seem to depend strongly on the resolution. The light lines show individual realisations of black holes at each mass ratio where only the dynamical friction correction is applied. In some cases, this faithfully recovers the black hole orbit ($M_{\rm BH}/M_{\rm DM}=2$, $2.5$, and $4.5$), but for others the black holes do not sink within the age of the Universe. Interestingly, this does not seem to depend strongly on the mass ratio in this regime (some black holes with $M_{\rm BH}/M_{\rm DM}=4$ and 5 do not sink even with a dynamical friction correction). We conclude that a stochasticity correction is required to ensure that a black hole sinks to the centre of a simulated halo, however the inherent randomness of this process makes it difficult to recover black hole orbits accurately for any individual case in the regime $M_{\rm BH}/M_{\rm DM}\leq 5$.
We have repeated this test with rescaling the dynamical mass of the black hole to $M_{\rm BH}+M_{\rm DM}$ instead of applying the stochasticity correction directly. We found that up to $M_{\rm BH}/M_{\rm DM}\leq 5$, 10 per-cent of the initial conditions lead to substantailly delayed sinking of the black holes (the black holes do not sink before 13 Gyr) and 13 per cent of initial conditions show no or little sign of sinking over 13 Gyr, while the median sinking times for each $M_{\rm BH}/M_{\rm DM}$ lie between 8 and 12 Gyr. Directly applying the stochastic correction, on the other hand, at all times guarantees the sinking of black holes.
@@ 5 Applying the model
Figure 11: Top: the median trajectories of $10^{9}\,{\rm M}_{\odot}$ black holes in a $10^{13}\,{\rm M}_{\odot}$ halo computed using our calibrated dynamical friction model. The results are shown for 7 mass resolutions and their corresponding optimal softening values (see Table 1). The grey dotted and dashed lines show the sinking time of the black hole at the highest background particle mass resolution using the Chen et al. (2022) and Tremmel et al. (2015) models, respectively. The arrows mark the earliest and the latest sinking time computed from 6 simulations at each resolution with differing random seeds in the initial conditions. The sinking time is defined as the time when the black hole first comes one softening length away from the centre of the halo. The left and right panels show the results with and without the application of the noise correction, respectively. Bottom: same as the top panel, but using a fixed softening of $\epsilon_{g}=0.44$ kpc.

### v2

@@ Abstract
pubyear: 2023pagerange: A calibrated model for N-body dynamical friction acting on supermassive black holesC
@@ 1.2 Compensating for unresolved dynamical friction
The method of Pfister et al. (2019) is similar to those described above, but takes into account the contribution of fast-moving particles to dynamical friction, which may be important in enhancing dynamical friction in cases where the density distribution of the background particles becomes shallow (Antonini & Merritt, 2012; Dosopoulou & Antonini, 2017), such as in central density cores (de Blok, 2010). However, we also note that the “standard” Chandrasekhar treatment is known to fail in reproducing the effect of stalling in cored density distributions (Read et al., 2006; Kaur & Sridhar, 2018; Banik & van den Bosch, 2021).
@@ 1.3 Dynamical friction from $N$-body particle distributions
Using their method, the dynamical friction force can be computed straightforwardly with direct summation, or with tree methods by treating a distant node as a single particle. Manifest momentum conservation can also be introduced by applying an opposite force to each particle that contributes dynamical friction to the black hole – a feature that cannot be self-consistently implemented alongside models of Tremmel et al. (2015) and Chen et al. 2022 (as well as the model presented in this work, see Appendix C). Nevertheless, for relatively well-resolved cases, such an approach, in effect, double-counts the dynamical friction because a fraction of the force is already resolved in the simulation and taken into account in the force computation. Ma et al. (2023) acknowledge this shortcoming and point out that this can be corrected in the future by some mass-resolution-dependent weight function.
@@ 2 Methods
Table 1: The black hole to dark matter mass ratio, the number of particles used, the number of particles within the original orbital radius of the black hole and the employed gravitational softening for our primary set of simulations carried out with GADGET-4. In bold, we show the resolution of the simulation carried out with KETJU.
@@
5000 $\mathbf{5\times 10^{7}}$ $\mathbf{2\times 10^{6}}$ 0.20
@@ 2.3 Predictions from simulations with KETJU
In order to make sure that the Chandrasekhar description works for our set-up, we also run a high-resolution $N$-body simulation using the KETJU code (Rantala et al., 2017; Mannerkoski et al., 2023). The initial conditions we use for this simulation are highlighted in bold in Table 1. The essence of the GADGET-4-based KETJU code is to add a regularized region around each massive black hole in the simulations in which the gravitational dynamics is accurately integrated without gravitational softening. Throughout this work we use the regularized region size of $2.8\,\epsilon_{g}$ in radius, resulting in region sizes up to a few hundreds of pc, ensuring that all interactions with the black hole in the simulations are always non-softened. The interactions between dark matter particles in the region remain softened. This allows us to capture the dynamical friction below the softening length, while also accounting for deviations from the Chandrasekhar approximation (e.g. like the assumption of hyperbolic orbits). Moreover, since a number of previous methods used to include unresolved dynamical friction in simulations rely on the Chandrasekhar description, we aim to learn about any potential shortcomings of these methods through this test.
In KETJU, the integration around black holes (the MSTAR integrator, Rantala et al. 2020; Mannerkoski et al. 2023) relies on algorithmic regularization techniques (Mikkola & Tanikawa, 1999; Mikkola & Merritt, 2006), a minimum spanning tree coordinate system and the Gragg-Bulirsch-Stoer (GBS) extrapolation method (Gragg, 1965; Bulirsch & Stoer, 1966) to achieve an extremely high integration accuracy. The efficient two-fold parallelization of MSTAR allows for integrating up to $\sim 5000$ particles in a single regularized region. In order to properly resolve dynamical friction, KETJU requires a mass ratio of $M_{\rm BH}/M_{\rm DM}\gtrsim 100$ between the black hole and the dark matter particles (Rantala et al., 2017). In this work, we use the mass ratio $M_{\rm BH}/M_{\rm DM}=5000$. KETJU has been successfully applied for simulating massive black hole and black hole binary dynamics in both isolated galaxy mergers (e.g. Rantala et al. 2018) and cosmological zoom-in simulations (Mannerkoski et al., 2022). Note that in this work we employ only Newtonian gravity and do not include post-Newtonian terms in the equations of motion close to the black hole, something that is in principle available in KETJU, as the dark matter mass elements do not represent physical particles. For the user-given accuracy parameters of KETJU, we adopt a GBS tolerance of $\epsilon_{\mathrm{GBS}}=10^{-7}$ and an integration end-time tolerance of $\epsilon_{\mathrm{t}}=10^{-2}$.
@@ 3 Performance of existing methods
The result of applying the Chen et al. (2022) formula can be seen in the middle panel of Fig. 2, where the black holes now seem to sink, on average, faster than the prediction of KETJU by up to 1 Gyr, and there appears to be a slight resolution-dependent gradient in the sinking time. This may suggest that $b_{\rm max}=6\,\epsilon_{g}$ in the Coulomb logarithm is too large of a dynamical friction correction, particularly in cases where dynamical friction is to a large part resolved.
@@ 4 A calibrated dynamical friction model
We note that the order of magnitude of the second-order coefficients is determined primarily by the mass of the background particles, while the first-order coefficient (the strength of dynamical friction) depends on the black hole mass. Thus, the overall effect of stochastic heating on the sinking timescale of the black hole depends sensitively on the ratio $M_{\rm BH}/M_{\rm DM}$. The coefficient in Eq. (11) acts opposite to the direction of the dynamical friction force and accounts for the velocity dispersion of the slow-moving background particles and the typical velocity of the fast-moving particles.
@@
Through explicit evaluation of Eqns. (10), (11) and (12), we have determined that for an isotropic Hernquist halo the second-order diffusion term acting parallel to the direction of motion of the black hole, ${\bf a}_{\rm noise,\parallel}$, approximately cancels out the additional cooling term, ${\bf a}_{\rm df,m_{*}}$, and only dominates near the centre of the halo. The second-order perpendicular term, ${\bf a}_{\rm noise,\perp}$, is approximately equal in magnitude to the parallel term, suggesting that Eq. (14) is a good approximation to the magnitude of the noise terms (see Appendix A). However, we also found that the Maxwellian approximation typically overestimates the stochastic term in a Hernquist halo by a factor of $\sim 3$, while the numerical evaluation of local $\sigma_{\rm 3D}/v$ is extremely noisy and can lead to order-of-magnitude changes in the black hole acceleration. Moreover, it is unclear how one would account for the second-order terms perpendicular to the black hole motion, as we cannot at any given point determine the net direction of the stochastic force.
@@ 4.3 Probing different values of the softening
$\eta_{\rm optimal}=\rm min[0.46~{}({\epsilon_{g}}/{\rm kpc})^{0.17}~{}10^{~{}0.11\ln^{2}(\epsilon_{g}/{\rm kpc})}~{},~{}1.0],$ (16)
where we fixed the maximum value of $\eta$ to 1. The relation $b_{\rm max}=\epsilon_{g}$ is shown with the grey dashed line. We note that the definition of $b_{\rm min}$ of Tremmel et al. (2015) and Chen et al. (2022) is different to ours, so our relation would not hold in that case; however, with $b_{\rm min}=GM/v^{2}_{\rm BH}$, and through an approximation $v^{2}_{\rm BH}\approx\frac{1}{3}\sigma_{\rm 3D}^{2}$, we can deduce that an optimal value is $\eta\approx 2$ for Tremmel et al. (2015) and Chen et al. (2022). We show the optimal value of $\eta$ with this definition of $b_{\rm min}$ in Appendix B.
@@
$\eta_{\rm optimal}=\rm min[103.7~{}({\epsilon_{g}/a_{\rm HQ}})^{1.3}~{}10^{~{}0.11\ln^{2}(\epsilon_{g}/a_{\rm HQ})}~{},~{}1.0].$ (17)
@@ 4.6 A treatment for low-mass black holes
Figure 9: Left: the effect of increasing the dynamical mass of black holes on their orbital trajectories. In black we show the analytical prediction for the orbit of a $10^{8}\,{\rm M}_{\odot}$ black hole in a $10^{13}\,{\rm M}_{\odot}$ halo, starting on a circular orbit at 10 kpc. The initial orbit with $M_{\rm BH}/M_{\rm DM}=1$ is shown with a thin dark blue line. The dynamical mass of the black hole is then scaled by a factor $k$ in the range 1 to 10 times the dark matter particle mass, resulting in faster black hole sinking. No dynamical friction correction is applied. The thick red line shows the result of setting the dynamical mass of the black hole to $M_{\rm dyn}=M_{\rm BH}+M_{\rm DM}$ and applying the correction for unresolved dynamical friction via Eqn. (8). Right: same as the left panel, but for a $10^{7}\,{\rm M}_{\odot}$ black hole starting at 5 kpc. The black hole trajectories have been smoothed over 0.1 Gyr for clarity.
@@
where $M_{\rm DM}$ is the background particle mass, which brings the noise correction to the right order of magnitude (see Appendix A). Since particle scattering below $\eta\epsilon_{g}$ is unresolved, only impact parameters above $\eta\epsilon_{g}$ contribute, giving the Coulomb logarithm $\ln\Lambda=\frac{R}{\eta\epsilon_{g}}$. We note that this is different to applying the term ${\bf a_{\rm df,m_{*}}}$ as in the work of Damiano et al. (2024) – their additional “cooling” force accounts for unresolved dynamical friction from the mass of background particles (assuming their masses are the true masses of background stars and dark matter), while Eq. (20) attempts to negate the resolved heating of the black hole by massive background particles.
@@
The result of the application of this stochasticity correction through a rescaling of the dynamical mass, together with the correction for unresolved dynamical friction, is shown with a thick red line in Fig 9. In neither case does the correction fully reproduce the true trajectory of the black hole, however, in both cases the black holes sink to the centre of the galaxy without the need for a substantial increase of their dynamical mass.
@@ 4.7 Black holes in the regime $1\leq M_{\rm BH}/M_{\rm DM}\leq 5$
Figure 10: The results of applying our dynamical friction correction, coupled with the correction for stochasticity, to a $10^{7}\,{\rm M}_{\odot}$ black hole in a $10^{13}\,{\rm M}_{\odot}$ halo starting at 4 kpc. Darker lines show median black hole trajectories for different $M_{\rm BH}/M_{\rm DM}$ ratios as the halo mass resolution changes. The error bars show the range of sinking times of black holes from the sample of 6 initial conditions used to compute the median trajectories. The lighter lines show examples of trajectories where only a dynamical friction correction is applied, i.e. no stochasticity correction. The black hole trajectories have been smoothed over 0.15 Gyr for clarity.
We now explore how our stochasticity correction behaves in the regime $1\leq M_{\rm BH}/M_{\rm DM}\leq 5$. This is shown in Fig. 10, where for a fixed halo and black hole mass, we gradually increase the mass resolution such that $M_{\rm BH}/M_{\rm DM}$ varies from 1 to 5. The dark lines show the median trajectory of 6 black hole seeds for each mass ratio. We apply the stochasticity correction via Eq. (20). The correction clearly tends, on average, to sink the black holes too fast, with the effect being strongest for lower $M_{\rm BH}/M_{\rm DM}$. The error bars range from the earliest to the latest sinking times of the black holes, where the mass ratio appears to set the earliest black hole sinking time, while the maximum sinking time does not seem to depend strongly on the resolution. The light lines show individual realisations of black holes at each mass ratio where only the dynamical friction correction is applied. In some cases, this faithfully recovers the black hole orbit ($M_{\rm BH}/M_{\rm DM}=2$, $2.5$, and $4.5$), but for others the black holes do not sink within the age of the Universe. Interestingly, this does not seem to depend strongly on the mass ratio in this regime (some black holes with $M_{\rm BH}/M_{\rm DM}=4$ and 5 do not sink even with a dynamical friction correction). We conclude that a stochasticity correction is required to ensure that a black hole sinks to the centre of a simulated halo, however the inherent randomness of this process makes it difficult to recover black hole orbits accurately for any individual case in the regime $M_{\rm BH}/M_{\rm DM}\leq 5$.
We have repeated this test with rescaling the dynamical mass of the black hole to $M_{\rm BH}+M_{\rm DM}$ instead of applying the stochasticity correction directly. We found that up to $M_{\rm BH}/M_{\rm DM}\leq 5$, 10 per-cent of the initial conditions lead to substantailly delayed sinking of the black holes (the black holes do not sink before 13 Gyr) and 13 per cent of initial conditions show no or little sign of sinking over 13 Gyr, while the median sinking times for each $M_{\rm BH}/M_{\rm DM}$ lie between 8 and 12 Gyr. Directly applying the stochastic correction via Eqn. (20), on the other hand, at all times guarantees the sinking of black holes.
@@ 5 Applying the model
Figure 11: Top: the median trajectories of $10^{9}\,{\rm M}_{\odot}$ black holes in a $10^{13}\,{\rm M}_{\odot}$ halo computed using our calibrated dynamical friction model. The results are shown for 7 mass resolutions and their corresponding optimal softening values (see Table 1). The grey dotted and dashed lines show the sinking time of the black hole at the highest background particle mass resolution using the Chen et al. (2022) and Tremmel et al. (2015) models, respectively. The error bars show the range of sinking times computed from 6 simulations at each resolution with differing random seeds in the initial conditions. The sinking time is defined as the time when the black hole first comes one softening length away from the centre of the halo. The left and right panels show the results with and without the application of the noise correction, respectively. Bottom: same as the top panel, but using a fixed softening of $\epsilon_{g}=0.44$ kpc.